7/6, 6/7, 8/9 from least to greatest please.
Found 2 solutions by jim_thompson5910, Alan3354:
Answer by jim_thompson5910(35256)

Use a calculator to get


7/6 = 1.166667
6/7 = 0.857143
8/9 = 0.888889

All decimal values above are approximate.

-------------------------------------------------------

So the list 7/6, 6/7, 8/9 is approximate to this list 1.166667, 0.857143, 0.888889 in the exact same order.

Now sort the list 1.166667, 0.857143, 0.888889 to get 0.857143, 0.888889, 1.166667

-------------------------------------------------------

We have this sorted list 0.857143, 0.888889, 1.166667

Replace that list with their fractional equivalent values to get this final answer: 6/7, 8/9, 7/6

Answer by Alan3354(69443)

You can find a common denominator, or use a calculator and compare the decimal values.
You're welcome.